### v2

@@ Methods
For all experiments, we set hyperparameters $\tau_{\mathrm{cyc}}=10$, $\tau_{\mathrm{PC}}=8$, and $\tau_{\mathrm{ESS}}=\tau_{\mathrm{idt}}=1$. We use Adam solver [69] to initiate training of the networks from scratch, maintaining a consistent learning rate of 0.0001 for the initial 50 epochs. Subsequently, we linearly decrease the learning rate to zero over the following 50 epochs. Each epoch processed 10,000 images (we set the input dimension as $256\times 256\,px$) with a batch size 20. The training was on a single Nvidia RTX A100 GPU; for dataset ‘GenLFI-Microfluidics’, it took 29 hours. All tests were conducted on a workstation (CPU Intel i9-13900K, GPU Nvidia RTX3090). We utilized FID for quantification and comparative evaluation. Following fast reconstruction, we performed post-processing for large sample results using image stitching algorithms. As illustrated in Fig.5b, we employed local object stitching [50] and global panorama stitching (see Supplementary Information 6) separately.
@@
Using the GenLFI setup, we recorded 500 hologram frames ($i$) in total (.fit format, resolution $6072\times 4042\,px$, acquired via the ASI Studio software) focusing on the ROI of the microfluidic chip with dynamic droplets (see Fig.4). Then, under the same experimental conditions, we repeated the experiment to acquire real-world bright field images ($o$) as the data guidance using a bright-field microscope. Due to the small FOV of the microscope, we scanned the same ROI and recorded a 5-minute video containing 7000 frames of bright-field images.
@@
Cell spheroids imaging
Human pleural mesothelial cells MeT-5A NF2KO were cultured as 3D spheroids by seeding approximately 30,000 cells into each well of a 6-well ultra-low attachment culture plate (#3471; Corning, Thermo Scientific); details see [70]. The dataset ‘GenLFI-Spheroids’ construction and network training for this task follows a similar routine of Microfluidics imaging. We used a red laser (see Optical setup) as the light source. The bright-field microscope for recording is Leica DMIRE2 with a 10x objective lens, coupling with a polarization filter to enhance the light contrast for spheroid structure observation. The acquired raw data contains 132 frames of holograms and 26,581 frames of bright-field microscopic images. We constructed a dataset comprising 17,418 training and 100 test images (captured from an unseen well) for LFI holograms (after Bicubic interpolation super-resolution, $808\times 808\,px$), and 23,738 training and 100 test images for bright-field images (original size, $1280\times 1024\,px$).
@@
We employed a USAF 1951 test chart (Thorlabs R3L3S1P) to evaluate the resolution limits of our LFI system. Given that the thickness of the test bench’s glass slide, i.e., $1.5\,mm$, affects the light diffraction, we placed the printed side close to the microscope’s lens to minimize this effect when capturing images through standard bright-field microscopy. Then, we placed the sample in the same orientation to acquire holograms with GenLFI. For microscopy images, we selected a square ROI with the smallest pair of lines. For holograms, we cropped out the same ROI to a $220\times 220\,px$ image (see Fig.2b). We carefully aligned these two domain images on the same scale. To augment our dataset for this task, we introduced Gaussian noise to generate 10,000 variant images for each domain, of which 99 images with 1 original noise-free image as test images, and the other images as training images.
@@
Multi-focus imaging
In evaluating multi-focus imaging capabilities, we stacked four glass slides, each $1\,mm$ thick, upon which handwritten numbers (1,2,3,4) had been inscribed. Another slack is marked with 5,6,7,8 at different heights as the unseen object for the test. These numbers serve as markers at different distances from the CMOS sensor’s imaging plane. When examined under bright-field microscopy, focusing on any one number caused the others to appear progressively more out of focus. We captured focused images of each number separately to establish clear ground truth data for training the LensGAN. The last dataset, ‘GenLFI-Multifocus’, consists of 43 test and 100 training images of hologram (1072×(669 or $1072\times(669\text{ or }1338)\,px$); and 27 test and 395 training images of bright-field images ($4096\times 2160\,px$). During training, hologram images were guided solely by these focused images.
@@ 1 Zebrafish imaging
    2 Further discussion on complex optical field
    Here we further discuss the optical field with large axial size samples and mathematically prove the inadequacy of conventional model-based 3D lens-free holographic microscopy (LFHM) tomographic reconstruction approaches. To this end, we analyze the current 3D LFHM algorithm based on scalar diffraction theory for 3D objects.
    As shown in Fig.1 in the main text, the solution to the light propagation problem is described as:
    $u(\boldsymbol{r})=u^{inc}(\boldsymbol{r})+u^{sct}(\boldsymbol{r})$ (11)
    where $u^{inc}(\boldsymbol{r})$ represents the incident light and $u^{sct}(\boldsymbol{r})$ represents the scattered light. The Lippmann-Schwinger equation (LSE) depicts the relationship between the 3D scattering potential of the sample and the resulting scalar field at region $V$:
@@ 2 Further discussion on complex optical field
    where $G(\boldsymbol{r})=\frac{1}{4\pi R}\exp(jk_{b}R)$, $k_{b}=k_{0}n_{b}$ is the wavenumber of the light in the background medium, and $R$ denotes the length of the Euclidean vector $\boldsymbol{r}$. Discretized and denoted by vectors, the resulting scalar field $\boldsymbol{\hat{u}}$:
    $\widehat{\boldsymbol{u}}=\widehat{\boldsymbol{u}}^{inc}+\widehat{\boldsymbol{G}}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}=\mathcal{L}(\boldsymbol{f})$ (13)
    where $\mathcal{L}(\cdot)$ is the light propagator, $\boldsymbol{\operatorname{diag}(f)}\in\mathbb{C}^{N\times N}$ denotes a diagonal matrix whose main diagonal entries consist of elements from the vector $\boldsymbol{f}$. With multiple light sources, the final forward model of the 3D LFHM can be presented by the intensity of the optical field $\widehat{\boldsymbol{H}}$ under $p^{th}$ illumination:
    $\widehat{\boldsymbol{H}}=\left|\widehat{\boldsymbol{u}}^{inc}+\widehat{\boldsymbol{G}}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}\right|^{2}=\left|\mathcal{L}_{p}(\boldsymbol{f})\right|^{2}$ (14)
@@
    $\boldsymbol{f}^{*}=\arg\min_{f\in\mathbb{C}^{N}}\{\Sigma_{p=1}^{P}d(\left|\mathcal{L}_{p}(\boldsymbol{f})\right|^{2},\widehat{\boldsymbol{H}}_{p})\},$ (15)
    where $\mathcal{L}_{p}(\cdot)$ denotes the light propagator, $\boldsymbol{f}^{*}$ denotes the amplitude of the light, $\widehat{\boldsymbol{H}}_{p}$ represents the measured light intensity under the $p^{th}$ illumination.
    In the microfluidics experiment, the mediums consist of air, PDMS, water, oil, and glass. In this condition, different components of the electromagnetic field may make the vector diffraction theory of $u(\boldsymbol{r})$ in (11) ineffective. Furthermore, the model-based algorithm’s approximation of the hologram in the 2D space is affected by the thickness of the sample. For example, the Born approximation for (13) is valid only if $\boldsymbol{u}^{sct}\ll\boldsymbol{u}^{inc}$. This corresponds to $l(n_{s}-n_{b})\ll\lambda$ , where $l$ is the thickness, $(n_{s}-n_{b})$ describes the medium’s refractive index variation, and $\lambda$ is the wavelength of the incident light. Consider a PDMS chip that has water inside, with a thickness of $5000\,\mu m$; there are $l(n_{s}-n_{b})$ = $5000\times(1.40-1.33)=350\,\mu m$, while $\lambda=0.42\,\mu m$. It is obvious that the model is not applicable to the condition of Born approximation for large samples with strong scattering.
    To provide a more accurate description of the physical process, some non-linear models have been proposed [30]. These iterative methods involve either multiple iterations by variations or dividing the sample into a series of thin layers. However, these methods are impractical to apply in real experiments and are restricted to sample-specific tools.
    In summary, real-world scattered fields are complex and difficult to model. The existing forward model-based LFI systems are limited to small and thin samples.
@@ 4 Discriminator design
    $\beta_{j,i}=\frac{\exp(s_{ij})}{\sum_{i=1}^{N}\exp(s_{ij})},\text{where }s_{ij}=f(x_{i}^{T})g(x_{j})$ (16)
    $\beta_{j,i}$ indicates the range of attention where the model computes mapping of the $i^{th}$ location of the $j^{th}$ feature regions. The output of the attention layer is $a=(a_{1},a_{2},...,a_{j},...,a_{N})\in\mathbb{R}^{C\times N}$ , where
    $a_{j}=v(\sum_{i=1}^{N}\beta_{j,i}h(x_{i})),h(x_{i}=W_{h}x_{i}),v(x_{i}=W_{v}x_{i})$ (17)
    $W_{g}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{f}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{h}\in\mathbb{R}^{\overline{C}\times C}$ , and $W_{v}\in\mathbb{R}^{C\times\overline{C}}$ are three learned weight metrics that are implemented as $1\times 1$ convolutions. Practically, for memory efficiency, the channel number of $\overline{C}$ is reduced to be $C/k$ , where $k$ is set as 8. By multiplying $a_{i}$ a learnable scale parameter $\gamma$ and adding back the input feature map $x_{i}$, the final scaled output is:
@@
    5 Ablation study
    Figure S4: Visual comparison of results by different I2I algorithms. We selected 3 cases from the ’Hologram2Microscopy’ dataset and evaluated different algorithms with FID: classical iterative holographic reconstruction algorithm [22], our algorithm LensGAN, CycleGAN [66], our backbone IrwGAN [63], IrwGAN enhanced with our self-attention discriminator, and SOTA unpaired I2I algorithm GP-UNIT [73].
    We evaluated LensGAN on an unseen test set (images from full ROI frames) with Fréchet Inception Distance (FID) [74], as depicted in Fig.S4. FID is the current standard for assessing the quality of generative models, where a lower FID indicates a better correlation between the distributions of generated images and ground truth. In the ablation study, we replaced the discriminator in IrwGAN [63] with our own (enhanced with self-attention blocks, details in Discriminator design), denoted as ‘IrwGAN + Self-attention.’ We also evaluated the CycleGAN [66] and a SOTA unpaired I2I model GP-UNIT [73] on our dataset.
    Fig.S4 shows the results that compared to the IrwGAN [63] backbone, the new discriminator improves the FID score by 3.27, and together with the $\mathcal{L}_{\mathrm{PC}}$, the LensGAN improves the FID score by 10.41. The results in Fig.S4 demonstrate that our model achieved the lowest score and successfully addressed the main distortions and artifacts. $\mathcal{L}_{\mathrm{PC}}$ constrains the generated results effectively, preventing the line artifacts in IrwGAN (see cases 1 & 3), and our discriminator $D$ improves the reconstructed droplets’ illumination details (see case 2).
@@ 8 Spheroid genotypes
    9 Failure case
    Figure S7: Failure case of singular sample: reconstruction for an organ-on-chip. We used a single raw hologram and BF microscopy image for training. Due to the low data scale, although the reconstruction achieved a high SSIM score, some information are lost in the reconstruction.

### v4

@@ Methods
The training and testing follow a standardized CycleGAN PyTorch routine [93], with details indicated in the supplementary code documentation. For all experiments, we set hyperparameters $\tau_{\mathrm{cyc}}=10$, $\tau_{\mathrm{PC}}=8$, and $\tau_{\mathrm{ESS}}=\tau_{\mathrm{idt}}=1$. We use Adam solver [94] to initiate training of the networks from scratch, maintaining a consistent learning rate of 0.0001 for the initial 50 epochs. Subsequently, we linearly decrease the learning rate to zero over the following 50 epochs. Each epoch processed 10,000 images (we set the input dimension as $256\times 256\,px$) with a batch size of 1 (to perform instance normalization). The training was on a single Nvidia RTX A100 GPU; for dataset ‘GenLFI-Microfluidics’, it took 29 hours. All tests were conducted on a workstation (CPU Intel i9-13900K, GPU Nvidia RTX3090). We utilized FID for quantification and comparative evaluation. Following fast reconstruction, we performed post-processing for large sample results using image stitching algorithms. As illustrated in Fig.4b, we employed local object stitching [71] and global panorama stitching (see Supplementary Information 5) separately.
Cell spheroids imaging
Human pleural mesothelial cells MeT-5A NF2KO were cultured as 3D spheroids by seeding approximately 30,000 cells into each well of a 6-well ultra-low attachment culture plate (#3471; Corning, Thermo Scientific); details see [95]. The dataset ‘GenLFI-Spheroids’ construction and network training for this task follows a similar routine of Microfluidics imaging. We used a red laser (see Optical setup) as the light source. The bright-field microscope for recording is Leica DMIRE2 with a 10x objective lens, coupling with a polarization filter to enhance the light contrast for spheroid structure observation. The acquired raw data contains 132 frames of holograms and 26,581 frames of bright-field microscopic images. We constructed a dataset comprising 17,418 training and 100 test images (captured from an unseen well) for LFI holograms (after Bicubic interpolation super-resolution, $808\times 808\,px$), and 23,738 training and 100 test images for bright-field images (original size, $1280\times 1024\,px$). The time-lapse growth experiment was recorded for 2 hours at room temperature, as the spheroids needed to be cultured at ${37}^{\circ}C$ in the incubator.
@@
Using the GenLFI setup, we recorded 500 hologram frames ($i$) in total (.fit format, resolution $6072\times 4042\,px$, acquired via the ASI Studio software) focusing on the ROI of the microfluidic chip with dynamic droplets (see Fig.5). Then, under the same experimental conditions, we repeated the experiment to acquire real-world bright-field images ($o$) as the data guidance using a bright-field microscope. Due to the small FOV of the microscope, we scanned the same ROI and recorded a 5-minute video containing 7000 frames of bright-field images.
@@
We employed a USAF 1951 test chart (Thorlabs R3L3S1P) to evaluate the resolution limits of our LFI system. Given that the thickness of the test bench’s glass slide, i.e., $1.5\,mm$, affects the light diffraction, we placed the printed side close to the microscope’s lens to minimize this effect when capturing images through standard bright-field microscopy. Then, we placed the sample in the same orientation to acquire holograms with GenLFI. For microscopy images, we selected a square ROI with the smallest pair of lines. For holograms, we cropped out the same ROI to a $220\times 220\,px$ image (see Fig.3a). We carefully aligned these two domain images on the same scale. To augment our dataset for this task, we introduced Gaussian noise to generate 10,000 variant images for each domain, of which 99 images with 1 original noise-free image as test images, and the other images as training images.
@@
Acknowledgement
R.L. was supported by the Global Joint PhD Studentship between KU Leuven and the University of Edinburgh. Y.Y. was supported by the European Research Council Starting Grant, Chancellor’s Fellowship, and Bayes Innovation Fellowship. C.H.: The early stages of this work were supported by the JHMRF and Worldwide Cancer Research (19-0238). K.P. is supported by an MRC Precision Medicine DTP Studentship. We thank Willemien Gosselé (KUL) and Mengguang Ye (UoE) for their assistance with the experiments.
Authors contributions
R.L. contributed to all facets of this study, encompassing conceptualization, algorithm design, experimentation, analysis, interpretation, figure creation, and manuscript writing. Y.Y. and X.C. shared supervisory responsibilities for this project. P.B. supervised the optics experiments and offered valuable domain expertise in optics. Z.L. contributed to the panorama stitching algorithm. Additionally, P.B., M.W., K.P., G.F., Y.F., and C.H. contributed by providing samples and expertise for experimentation.
Competing interests
The authors declare no competing interests.
Data and code availability
Data and codes for the implementation are available at https://github.com/RL-arch/LensGAN.
@@ 1 Zebrafish imaging
2 Further discussion on complex optical field and 3D LFI
To image 3D samples, modern digital in-line holography systems often employ measurement diversity strategies, as shown in Fig.S2. These include capturing multiple holograms by varying parameters such as sample-to-sensor distances [96] and wavelengths [97] in 2D LFI or illumination angles [67] in 3D LFI. Furthermore, achieving high-resolution holograms frequently further relies on sub-pixel super-resolution via lateral (horizontal) shifting [19, 67]. Especially, lens-free 3D tomography (3D LFI, see Fig.S2c) is a 3D strategy for solving complex optical fields through multi-angle illumination.
Figure S2: Common multi-measurement strategies for LFI. a Multi-height phase recovery (MHPR) [96]: captures holograms at several sample-to-sensor ($z_{2}$) distances. b Multi-wavelength retrieval [97]: uses sequential illumination at five distinct wavelengths. c Lens-free 3D tomography [98]: records holograms under varied illumination angles for volumetric reconstruction.
Scalar diffraction theory
One current solution for volumetric samples is 3D tomographic lens-free imaging (3D LFI), as illustrated in Fig.S2c. However, these methods, like FBP (filtered back-projection) [98] or ODT (optical diffraction tomography) [99], also rely on the forward models. Here we discuss the optical field with large axial size samples and mathematically demonstrate the limitations of conventional model-based 3D lens-free holographic microscopy (LFHM) tomographic reconstruction approaches. To this end, we analyze the current 3D LFHM algorithm based on scalar diffraction theory for 3D objects.
Given a spatial vector $\mathbf{(}x$, the unique solution to the light propagation problem is described as [100]:
$u(\mathbf{x})=u^{inc}(\mathbf{x})+u^{sct}(\mathbf{x})$ (11)
where $u^{inc}(\mathbf{x})$ represents the incident waves and $u^{sct}(\mathbf{x})$ represents the scattered waves, as shown in Fig.1b. The Lippmann-Schwinger equation (LSE) depicts the relationship between the 3D scattering potential of the sample and the resulting scalar field at region $V(\mathbf{x})$:
@@ 2 Further discussion on complex optical field and 3D LFI
where $G(\mathbf{x})=\frac{1}{4\pi R}\exp(jk_{b}R)$ denotes Green’s function for the unperturbed wave equation, $k_{b}=k_{0}n_{b}$ is the wavenumber of the light in the background medium, $f(\mathbf{x})=k_{0}^{2}(n^{2}(\mathbf{x}-n_{b}^{2})$ defines the scattering potential, and $R$ denotes the length of the Euclidean vector $\mathbf{x}$. With all variants discretized and denoted by vectors, the resulting scalar field $\boldsymbol{u}$ can be represented by:
$\boldsymbol{u}=\boldsymbol{u}^{inc}+\boldsymbol{G}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}=\mathcal{P}(\boldsymbol{f})$ (13)
where $\mathcal{P}(\cdot)$ is a light propagator, $\boldsymbol{\operatorname{diag}(f)}\in\mathbb{C}^{N\times N}$ denotes a diagonal matrix whose main diagonal entries consist of elements from the vector $\boldsymbol{f}$. Here, $\mathcal{P}(\boldsymbol{f})$ describes the optical filed crafted.
In LFI, only the intensity (amplitude) of $\mathcal{P}(\boldsymbol{f})$, i.e., $\left|\mathcal{P}(\boldsymbol{f})\right|^{2}$, from Eq.(13) is measured. With multiple light sources, the final forward model of the 3D LFHM can be presented by the intensity of $\boldsymbol{H}$ under $k^{th}$ illumination:
$\boldsymbol{H}_{k}=\left|\boldsymbol{u}^{inc}+\boldsymbol{G}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}\right|^{2}=\left|\mathcal{P}_{k}(\boldsymbol{f})\right|^{2}$ (14)
@@
$\boldsymbol{f}^{*}=\arg\min_{f\in\mathbb{C}^{N}}\{\Sigma_{k=1}^{n}d(\left|\mathcal{P}_{k}(\boldsymbol{f})\right|^{2},\boldsymbol{H}_{k})\},$ (15)
where $\mathcal{P}_{k}(\cdot)$ denotes the light propagator, $\boldsymbol{f}^{*}$ denotes the amplitude of the light, which is the measured resolvable hologram.
Boundary condition
In the microfluidics experiment in Section Microfluidics screening, the mediums consist of air, PDMS, water, oil, and glass. In this condition, different components of the electromagnetic field may make the vector diffraction theory of $u(\mathbf{x})$ in Eq.(11) ineffective. Furthermore, the model-based algorithm’s approximation of the hologram in the 2D space is affected by the thickness of the sample. For example, due to the boundary condition of the continuity of the wave function and its derivative, the Born approximation for Eq.(13) is valid only if $\boldsymbol{u}^{sct}\ll\boldsymbol{u}^{inc}$. This corresponds to $l(n_{s}-n_{b})\ll\lambda$ , where $l$ is the sample thickness, $(n_{s}-n_{b})$ describes the medium’s refractive index variation, and $\lambda$ is the wavelength of the incident light. Consider a PDMS chip that has water inside, with a thickness of $5000\,\mu m$; there are $l(n_{s}-n_{b})$ = $5000\times(1.40-1.33)=350\,\mu m$, while $\lambda=0.42\,\mu m$. It should become obvious that the model does not apply to the boundary condition of Born approximation for large samples with intensive scattering.
To accurately describe the physical process, some non-linear models have been proposed [36]. These iterative methods involve either multiple iterations by variations or dividing the sample into a series of thin layers. However, these methods are impractical to apply in real experiments and are restricted to sample-specific tools.
In summary, for dynamic 3D samples, the real-world scattered fields are complex and challenging to model due to real-time variations in sample geometry and refractive index. Current LFI tomography is therefore limited to static, and relatively thin samples.
@@ 4 Discriminator design
$\beta_{j,i}=\frac{\exp(s_{ij})}{\sum_{i=1}^{N}\exp(s_{ij})},\text{where }s_{ij}=f(x_{i})^{T}g(x_{j})$ (16)
$\beta_{j,i}$ indicates the range of attention where the model computes mapping of the $i^{th}$ location of the $j^{th}$ feature regions. The output of the attention layer is feature map $a=(a_{1},a_{2},...,a_{j},...,a_{N})\in\mathbb{R}^{C\times N}$ , where
$a_{j}=v(\sum_{i=1}^{N}\beta_{j,i}h(x_{i})),$ (17)
$h(x_{i})=W_{h}x_{i},$
$v(x_{i})=W_{v}x_{i}$
$W_{g}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{f}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{h}\in\mathbb{R}^{\overline{C}\times C}$ , and $W_{v}\in\mathbb{R}^{C\times\overline{C}}$ are three learned weight matrices that are implemented by $1\times 1$ convolutions. Practically, for memory efficiency, the channel number of $\overline{C}$ is reduced to be $C/k$ , where $k$ is set as 8. By multiplying $a_{i}$ a learnable scale parameter $\gamma$ and adding back the input feature map $x_{i}$, the final scaled output is:
@@ 7 Spheroid genotypes
8 FOV comparison
The FOV comparison of different imaging systems is shown in Fig.S7.
Figure S7: FOV comparison, showing different imaging systems’ FOV on our microfluidic chip design. 4x: 4x objective of the bright-field microscope; state-of-the-art LFI [75]: the existing LFI system’s one-shot FOV based on holographic in-line reconstruction. AX 1x: Nikon AX [104], the world’s largest commercial confocal microscope. The scanning 1x objective can cross a $25\,mm$ diagonal FOV in multiple shots. With the low-resolution telescope setup, our FOV can reach $>500\,mm^{2}$.
9 Shape Preservation Score (SPS) Evaluation
To assess structural consistency between segmented objects in hologram and bright-field (BF) images from aligned datasets, we introduce the Shape Preservation Score (SPS). SPS quantifies the similarity of geometric shape descriptors between the two domains, measuring how well object-level morphological features are preserved during reconstruction.
Objective
The goal of SPS is to evaluate the fidelity of shape reconstruction by comparing statistical distributions of key geometric descriptors computed from objects segmented in the hologram and BF domains.
Shape Descriptors
We use three standard geometric descriptors [55] to characterize object morphology:
  1. 1.
    Aspect Ratio (AR): Measures the elongation of an object: $\text{AR}=\frac{L_{\text{major}}}{L_{\text{minor}}+\epsilon}$ where $L_{\text{major}}$ and $L_{\text{minor}}$ are the lengths of the major and minor axes of the ellipse fitted to the object. $\epsilon$ is a small positive constant (e.g., $10^{-6}$) added to prevent division by zero.
  2. 2.
    Circularity (Circ): Reflects how close an object’s shape is to a perfect circle:
    $\text{Circ}=\frac{4\pi A}{P^{2}+\epsilon}$
    where $A$ is the object’s area and $P$ its perimeter. Higher values indicate rounder shapes.
  3. 3.
    Solidity (Sol): Quantifies the convexity of the object:
    $\text{Sol}=\frac{A}{A_{\text{convex hull}}}$
    where $A_{\text{convex hull}}$ is the area of the convex hull enclosing the object. Solidity equals 1 for perfectly convex shapes.
Metric Definition
For each descriptor $d\in\{1,2,3\}$, we extract the empirical distributions $p_{d}(\text{holo})$ and $p_{d}(\text{bf})$ from all segmented objects in the hologram and BF images, respectively. We then compute the discrepancy between these distributions using the first-order Wasserstein distance ($W_{1}$, also known as Earth Mover’s Distance) [56].
The overall SPS is defined as:
$\text{SPS}=\frac{1}{D}\sum_{d=1}^{D}W_{1}\left(p_{d}(\text{holo}),\ p_{d}(\text{bf})\right)$
where $D=3$ is the number of descriptors.
Interpretation
SPS provides a scalar measure of morphological consistency across imaging modalities. A lower SPS value indicates closer alignment of object shapes between the hologram and BF domains, implying better structural preservation. In contrast, a higher SPS suggests that shape features have been distorted during reconstruction, leading to domain mismatch.
10 Fourier Ring Correlation (FRC) evaluation
Fourier Ring Correlation (FRC) has become a popular and relatively unbiased method of estimating image quality/resolution recently, especially in localization-based superresolution microscopy [105, 106]. FRC quantifies the correlation of spatial frequency content—specifically the magnitude of Fourier coefficients—between two images across concentric frequency bands, providing an objective measure of structural agreement at different resolutions.
  • Let $\mathcal{F}_{1}(u,v)$, $\mathcal{F}_{2}(u,v)$: Fourier transforms of the two images (shifted to center);
  • Let $r(u,v)=\sqrt{(u-u_{0})^{2}+(v-v_{0})^{2}}$ be the radial distance from center;
  • For each radius $r_{i}$, define a ring $R_{i}$ (pixels where $r(u,v)=r_{i}$);
Then FRC at ring $i$ is:
$\text{FRC}(r_{i})=\frac{\sum_{(u,v)\in R_{i}}|\mathcal{F}_{1}(u,v)||\mathcal{F}_{2}(u,v)|}{\sqrt{\sum_{(u,v)\in R_{i}}|\mathcal{F}_{1}(u,v)|^{2}\cdot\sum_{(u,v)\in R_{i}}|\mathcal{F}_{2}(u,v)|^{2}}+\epsilon}$
The overall FRC score is:
$\text{FRC}_{\text{mean}}=\frac{1}{N}\sum_{i=1}^{N}\text{FRC}(r_{i}),$
where $N$ is the number of radial frequency bins (rings), $\epsilon$ denotes a small constant to prevent division by zero;
A higher FRC means better preservation of frequency content; A score near 1.0 indicates a near-perfect correlation in Fourier magnitude structure.
We use FRC to evaluate the impact of the phase-contrast loss $\mathcal{L}_{\text{pc}}$. As shown in Fig. S8, three ROIs from Fig. 2 in the main text were analyzed in an ablation study. With $\mathcal{L}_{\text{pc}}$, LensGAN consistently yields higher FRC scores across all ROIs, reflecting improved frequency-domain agreement and finer structural fidelity. This demonstrates that phase-contrast loss enhances the preservation of high-frequency spatial features, much like phase-contrast microscopy, which reveals subtle structural variations in transparent samples.
Figure S8: Supplementary to Fig 2, ablation study on phase-contrast loss evaluated by FRC. Scale bar: 100 $\mu m$.
11 Performance under noise conditions
To evaluate the robustness of the model under varying noise conditions, we assessed LensGAN’s performance under different noise levels. Specifically, with the same pretrained LensGAN model, we introduced additive Gaussian noise to the clean input images of the paired dataset [59].
Given an original image $I$, the noisy image $I_{\text{noisy}}$ is generated as:
$I_{\text{noisy}}=I+\mathcal{N}(0,\sigma^{2})$
where $\mathcal{N}(0,\sigma^{2})$ denotes Gaussian noise with zero mean and standard deviation $\sigma$. We tested noise levels at $\sigma\in\{2,4,6,8,10,15,20,30,40,50\}$ to simulate a range of low to high noise conditions. The noise is applied pixel-wise, and values are clipped to the valid image intensity range (e.g., [0, 255] for 8-bit images). The results are illustrated in Fig.S9.
Figure S9: Performance under varying noise conditions are analyzed. The chart illustrates the lowest, highest, and mean average SSIM from 932 test sets.
Our results demonstrate that image quality degrades consistently with increasing Gaussian noise levels. The SSIM mean decreased from 0.7433 at $\sigma=0$ to 0.4046 at $\sigma=50$, indicating a substantial loss of structural similarity under high noise of $\sigma>10$. It is worth noting that the model was trained solely on clean, noise-free data; in practice, incorporating noise during training could substantially improve robustness to real-world imaging conditions.
12 Multi-focus experiment
Unlike conventional light microscopes requiring manual focus adjustments, GenLFI excels at reconstructing in-focus images from holograms containing objects at different depths. This unique capability is attributed to its virtual depth of field, offering significant flexibility for multi-focal imaging. To showcase this advantage, we conducted experiments with two stacks of 4 glass slides ($1\,mm$ thick), each marked with a written digit (font-weight $0.3\,mm$). In traditional bright-field setups, focusing is limited to a single depth plane, resulting in only one clear digit at a time (see Fig.S10). With proper data guidance, LensGAN can refocus and resolve all digits across different slides in a single reconstruction, achieving a superior depth of focus of 3 mm. This indicates our system’s versatility and underscores its potential in scenarios where simultaneous multi-depth clarity is crucial.
Figure S10: Results of handwritten digits slides stack. Left shows the 4x objective focuses on “1” and is gradually out of focus from “2” to “4”. Right is the reconstructed result; all numbers are focused.